OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 3x. What is the value of (g o f)(3)?

OpenStudy (princeharryyy):

gf(3) = 3(-2*3 + 8) = 3*(2) = 6

OpenStudy (princeharryyy):

@Student9623

OpenStudy (anonymous):

Where did you get the second 3 from the inside?

OpenStudy (anonymous):

Also what could be a possible decomposition of this be?

OpenStudy (princeharryyy):

gf(x) = 3(-2x + 8)

OpenStudy (anonymous):

Ohh okay that makes more sense thanks.

OpenStudy (princeharryyy):

and u need to find gf(3) so put 3 in the value of x

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

Could you look at the other one I posted for me because I cant figure out the decomposition of h(x).

OpenStudy (princeharryyy):

what do u want to kno in this problem

OpenStudy (princeharryyy):

f(x) or g(x) @Student9623

OpenStudy (anonymous):

I just need to know how to break it up into the inner and outer parts.

OpenStudy (princeharryyy):

see u kno if you are given gof(x) it means the function f(x) has been put inside g(x) that means it is actually g(f(x)) so here f(x) = x +3 and g(x) = 1/(x) ^2 @Student9623

OpenStudy (anonymous):

What? I was talking about this picture.

OpenStudy (princeharryyy):

yup this is for wat I gave u the answer @Student9623

OpenStudy (princeharryyy):

f(x) = x +3 and g(x) = 1/(x) ^2 if you put f(x) in g(x) u will get h(x) @Student9623

OpenStudy (anonymous):

Ohh sorry I was reading that wrong.

OpenStudy (princeharryyy):

wat did u read @Student9623

OpenStudy (anonymous):

I thought you were talking about the other one. Could you help me with this question also? Which of the following is the equation of a line parallel to the line y = 3x + 2, passing through the point (10,1)?

OpenStudy (anonymous):

I got 2x-3y=21

OpenStudy (princeharryyy):

ohk !! were u able to get the earlier one

OpenStudy (anonymous):

yes

OpenStudy (princeharryyy):

let the equation be y = 3x+c where c is some constant . now put 1 in place of y and 10 in place of x

OpenStudy (princeharryyy):

so, 1 = 3*10 +c => c = -29 , So, ans => y = 3x -29

OpenStudy (princeharryyy):

@Student9623

OpenStudy (princeharryyy):

parallel equations have same slope @Student9623

OpenStudy (princeharryyy):

has*

OpenStudy (anonymous):

ohh okay. Also would 4x-y+16 and 2x=3=-18 in matrix form be [4 -1] [2 3]?

OpenStudy (anonymous):

4x-y=16 ** sorry

OpenStudy (anonymous):

and 2x+3=-18 my bad I kept mixing up the keys

OpenStudy (princeharryyy):

2x=3=-18 wats this

OpenStudy (princeharryyy):

wouldn't it be 2x + 3y = 18

OpenStudy (anonymous):

I mean to put a plus sign in the middle and accidental put a equal sign also thats a negative 18

OpenStudy (princeharryyy):

so there isn't a y multiplied in 3 I was asking that @Student9623

OpenStudy (anonymous):

I need that in matrix form

OpenStudy (anonymous):

and i think its this

OpenStudy (princeharryyy):